OpenStudy (anonymous):

A circular pillar candle is 2.8 inches wide & 6 inches tall. What are the lateral area & surface area of the candle?

OpenStudy (lgbasallote):

the formula for lateral surface area is \[\pi d h\] does that help?

OpenStudy (lgbasallote):

d = diameter = wide h = height = tall

OpenStudy (anonymous):

so its the same for the surface area too?

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

surface area is \[2\pi r^2 + \pi d h\] r is radius

OpenStudy (anonymous):

so i would get 52.8 and 58.9

OpenStudy (lgbasallote):

is that what the calculator said?

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

then that's teh right answer...i have no calculator to verify sorry